Fermi National Laboratory


Accelerator Update for April 8, 2005 - April 11, 2005

Friday April 8
The day shift began with Operations monitoring store 4081 and the Recycler stash, with the Antiproton Source (Pbar) stacking, and with MiniBooNE and Meson MCenter experiments taking beam.

At 9:05 AM, Operations reported that beam was not making it into the Accumulator. An expert investigated, found that the problem was with a kicker (A:IKIK), and fixed it by raising the Thyratron gas pressure.

Operations terminated store 4081 at 10 AM. The Run Coordinator allowed the following maintenance work: Tevatron (TeV) sector D3 wet engine flywheel change, D3 QPM work, A3 power supply work, TeV RF strainer maintenance, TeV BPM maintenance, and to allow the Collider Detector at Fermilab (CDF) to access their collision hall.

The D-Zero (D0) experiment decided they would access their collision hall at 10:45 AM.

Operations established beam to MCenter at 10:56 AM.

Pbar began a stacking study at 12 PM.

Both D0 and CDF were out of access by 12:06 PM.

Although the A3 power supply work was the one job still underway at 12:46 PM, Operations began the TeV checkout and turn on.

Technicians reported at 1:41 PM that they had finished their A3 power supply work.

The Pbar studies ended at 1:50 PM.

The Neutron Therapy Facility (NTF) requested beam for calibrations at 2:32 PM.

Operations put the TeV through a dry squeeze at 3:27 PM.

A TeV expert reported at 3:44 PM that there might be a problem with the Front End for the A0 collimator.

Operations put the TeV through a wet squeeze at 4:02 PM.

Operations set the TeV to a 150 GeV ramp to test the A0 collimators at 5:10 PM.

Operations put the Recycler and the TeV into shot setup to establish a Recycler only store at 5:48 PM.

At 7:49 PM, shortly after all the antiprotons had been transferred into the TeV, the TeV suffered a sector A1 quench due to a separator spark.

Pbar resumed stacking at 7:53 PM.

MCenter resumed taking beam at 8 PM.

At 9:15 PM, a Booster RF station (BRF16) tripped off. It reset okay.

At 11:05 PM, the Pbar stack rate dropped very low. Operators investigated.

Saturday April 9
The midnight shift began Pbar stacking, and with MiniBooNE and MCenter taking beam.

Operations put the TeV into shot setup at 12:24 AM.

At 1:36 AM, some kind of a problem developed with the Main Injector (MI). At first it was thought to be a coalescing problem, but it wasn't. An operator called in an expert.

The MI expert reported at 3:05 AM that a RF cavity had a LCW flow switch that was stuck. Operators beat the flow switch until it gave up and resumed operating.

Operations established store 4088 at 5:37 AM with an initial luminosity of 75.04E30.

Operations tried to get Pbar to resume stacking at 5:38 AM, but it tripped off due to a problem with lithium lens power supply. Experts were called in.

MCenter resumed taking beam at 6 AM.

At 6:33 AM, a Linac RF station (KRF5) tripped off for the first of three times in five minutes. It reset okay each time.

At 6:48 AM, BRF16 tripped off at 6:48 AM.

At 10:18 AM, BRF16 tripped off for the first of three times in three hours.

The Pbar expert reported that his lithium lens power supply work was finished. Pbar resumed stacking at 12:49 PM.

At 1:45 PM, BRF4 tripped off for the first of many times in a short amount of time. The stations was bypassed.

At 2:14 PM, BRF16 tripped off. It reset okay.

At 4:22 PM, BRF6 tripped off. It took a while, but it was eventually reset.

At 4:54 PM, BRF6 tripped off. It reset okay.

Operations put the Recycler into shot setup to stash antiprotons at 6:16 PM.

At 6:29 PM, two Booster vacuum valves tripped off, which in turn tripped off GMPS and the chopper. Operators opened the valves and turned everything back on.

Operations reported at 7:04 PM that the Recycler stash had completed and Pbar had resumed stacking at 7:16 PM.

At 7:48 PM, BRF6 tripped off. Operations called in an expert to investigate.

Sunday April 17
The midnight shift began with Operations monitoring store 4088 and the stash, and with MiniBooNE and MCenter taking beam.

Operations terminated store 4088 at 8 AM.

Operations put the TeV into shot setup at 8:45 AM.

At 9:03 AM, Operations reported a problem with the TeV BPMs not saving data. Operators could not repair the problem, so they decided to continue with the TeV shot setup. Operations contacted an expert.

At 10:42 AM, at an expert's request, operators disconnected BRF6 from its power supply (EAPS).

Operations established store 4089 at 11:01 AM with an initial luminosity of 101.4E30. Pbar resumed stacking.

MCenter resumed taking beam 11:19 AM.

At 2:09 PM, a TeV vertical separator sparked three times. The sparking didn't seem to disturb the beam.

At 2:18 PM, a Booster power supply (EAPS) tripped off. Operators investigated and found its breaker tripped off. It would not stay reset. Operators called an expert for help and following his instructions, cleared the faults and turned the supply back on.

Pbar resumed stacking at 2:41 PM.

At 2:54 PM, the Booster EAPS tripped off and wouldn't reset. Operators called in an expert. At 3:28 PM, the expert reported that problem was due to BRF6, which he resolved. By 3:53 PM, Pbar had resumed stacking and MCenter resumed taking beam.

At 5:02 PM, a Linac quadrupole power supply (QPS317) had its current drop by twenty amps. Operators raised the current back up, but five minutes later the supply became very noisy. Operators swapped out the appropriate control cards and this solved the problem.

Operations put the Recycler into shot setup at 7:26 PM.

The Recycler stash ended and Pbar resumed stacking at 9:29 PM. MCenter resumed taking beam a few minutes later.

Monday April 11
The midnight shift began with operations monitoring store 4089 and the stash, with Pbar stacking and with MiniBooNE and MCenter taking beam.

At 1:27 AM, two Booster vacuum valves closed. Operators investigated but saw no problem. They opened the valves and turned everything back on.

At 2:11 AM, two Booster RF stations (BRF7 & 8) tripped off. They both reset okay.

The Plan for Monday
The plan for today is to terminate store 4089 around 2 PM, allow the experiments a two-hour access, and then put a combination store into the TeV.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 4/11/2005   email Fermilab